OpenStudy (anonymous):

I need to write an equation in slope intercept form for the line that satisfies each set o conditions..... Slope -2, passes through (-4,6)

OpenStudy (anonymous):

listen to me grasshopper : what you want is this: y-y1=m(x-x1) now (x1,y1) is your given point and m is your given slop. Plug then in and you will get your answer fo sho

OpenStudy (anonymous):

which would be -y1=m(x-x1)-y

OpenStudy (anonymous):

-y1=2(x-6)+4

OpenStudy (anonymous):

-y1=2x-8

OpenStudy (anonymous):

y=-2x+8...right?

OpenStudy (anonymous):

no son, what me :